Coldstream Corp. is comparing two different capital structures. Plan I would result in 13,000 shares of stock and $100,000 in debt. Plan II would result in 10,500 shares of stock and $150,000 in debt. The interest rate on the debt is 10 percent.

a. Ignoring taxes, compare both of these plans to an all-equity plan assuming that EBIT will be $90,000. The all-equity plan would result in 18,000 shares of stock outstanding. What is the EPS for each of these plans? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All equity $


b.
In part (a), what are the break-even levels of EBIT for each plan as compared to that for an all-equity plan? (Do not round intermediate calculations.)

EBIT
Plan I and all-equity $
Plan II and all-equity $


c.
Ignoring taxes, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations.)

EBIT           $  

d-1
Assuming that the corporate tax rate is 40 percent, what is the EPS of the firm? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

EPS
Plan I $
Plan II $
All equity $


d-2
Assuming that the corporate tax rate is 40 percent, what are the break-even levels of EBIT for each plan as compared to that for an all-equity plan? (Do not round intermediate calculations.)

EBIT
Plan I and all-equity $
Plan II and all-equity $


d-3
Assuming that the corporate tax rate is 40 percent, at what level of EBIT will EPS be identical for Plans I and II? (Do not round intermediate calculations.)

EBIT           $ ______
Answer #1

a)

EPS
PLAN 1 $ 6.15
PLAN 2 $ 7.14
All equity $ 5.00

Note : EPS = Earning attributable to equity shareholders / No. of shares

b) At the break even level, both EPSs should be equal. Let us equate them and solve for EBIT -

Plan 1 and all equity

EBIT/18000= EBIT-10000/13000

EBIT = 18000(EBIT-10000)/13000

EBIT = $36000

Plan 2 and all equity

EBIT/18000= EBIT-15000/10500

EBIT = $36000

EBIT
Plan 1 and all equity $ 36000
Plan 2 and all equity $ 36000

C) Plan 1 and Plan 2

EBIT-10000/13000= EBIT-15000/10500

EBIT = $36000

d-1)

EPS
PLAN 1 $ 3.69
PLAN 2 $ 4.28
All equity $ 5.00

Note : EPS = Earning attributable to equity shareholders / No. of shares

d-2)

At the break even level, both EPSs should be equal. Let us equate them and solve for EBIT -

Plan 1 and all equity

EBIT*(1-Tax)/18000= (EBIT-10000)(1-Tax)/13000

EBIT*(1-.40)/18000= (EBIT-10000)(1-.40)/13000

EBIT = $36000

Plan 2 and all equity

EBIT*(1-Tax)/18000= (EBIT-15000)(1-Tax)/10500

   EBIT*(1-.40)/18000= (EBIT-15000)(1-.40)/10500

EBIT = $36000

EBIT
Plan 1 and all equity $ 36000
Plan 2 and all equity $ 36000

d-3)

Plan 1 and Plan 2

(EBIT-10000)(1-Tax)/13000= (EBIT-15000)(1-Tax)/10500

(EBIT-10000)(1-.40)/13000= (EBIT-15000)(1-.40)/10500

  EBIT = $36000